# The large scale geometry of inverse semigroups and their maximal group images

### v1

@@ 4. Distortion and (weak) $F$-Inverse properties
    To prove that $S/\sigma=G$, observe that the group defined by regarding the presentation above as a group presentation is indeed $G$, as (R1) ensures $[e]_{G}=1$, (R2) yields a presentation for $G$ by definition, and the relations in (R3) are then automatic. The fact that $S$ is $E$-unitary follows immediately from the relations in (R1) and (R2).
    Now let $\Gamma$ denote the Cayley graph of $G$ with respect to the generating set $X\cup\{e\}$. Clearly $\Gamma$ is obtained from the Cayley graph of $G$ with respect to $X$ by adding an $e$-labelled loop at each vertex. Let $\Omega$ be the subgraph of $\Gamma$ consisting of $\Delta$ with an extra $e$-labelled loop at $1$. We claim that $\Omega$ satisfies conditions of Lemma 2.3 to be isomorphic to the Schützenberger graph of $e$. Indeed, it clearly contains the path (in other words, the loop) labelled $e$ starting at $1$. It remains to show that if one side of a defining relation of $S$ labels a path between two vertices in $\Omega$, then the other side of the relation labels a path between the same two vertices. For the relations of type (R1) and (R2) notice that if $w=e$ or $w\in(X\cup X^{-1})^{\ast}$ with $[w]_{G}=1$ then every path in $\Gamma$ labelled $w$ is closed; clearly this implies that $w$ labels a path between two vertices if and only if $w^{2}$ does so.
@@
The inverse monoid $S=\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}c^{-1}cd^{-1}d=1\rangle$ is an $E$-unitary inverse monoid without locally bounded group distortion.
@@
First, observe that the inverse monoid is the same as the one defined by the presentation $\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1,c^{-1}c=1,d^{-1}d=1\rangle$. This is an idempotent pure quotient of the one-relator inverse monoid $\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1\rangle$, which is $E$-unitary by [14] since the relator is a cyclically reduced word. Any idempotent pure quotient of an $E$-unitary inverse semigroup is $E$-unitary [16, Section 7.5, Proposition 4].
The maximal group image is $G=\operatorname{Gp}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1\rangle$. Let us add a new generator $u=bcb^{-1}$; then
@@
For convenience, we will identify the elements of $G$ with elements of $\operatorname{FG}(a,b,u)$, that is, with freely reduced words. It is helpful to keep this picture in mind depicting the relations in $G$:
$c$$d$$u$$a$$b$$a$$b$
The Cayley graph $\Gamma(G)$ with respect to the generating set $a,b,c,d$ is obtained from the Cayley graph of $\operatorname{FG}(a,b,u)$ (with respect to the generating set $a,b,u$) by replacing every $u$-edge with the surrounding rectangle and determinizing. In particular, observe that the black edges of the graph below form a subgraph of $\Gamma(G)$:
@@
$c$$d$$a$$b$$a$$b$$c$$d$
In particular, $S\Gamma(w)$ contains the graph below at every vertex $v$.
@@
We furthermore claim that there exists some vertex $v$ of $S\Gamma(w)$ such that no path of the form $bc^{-k}b^{-1}$, or $ad^{-k}a^{-1}$ $(k\in\mathbb{N})$ is readable from $v$ in $S\Gamma(w)$. Note that $[bc^{-k}b^{-1}]_{G}=[ad^{-k}a^{-1}]_{G}=u^{-k}$, so it suffices to show that there exists some vertex $v$ of $\Gamma$ such that $vu^{-k}\notin V(S\Gamma(w))$ for any positive $k$.
Let $w_{1},\ldots,w_{n}\in G$ be the vertices traversed by $w$. Then for any $v\in V(S\Gamma(w))$, there is a sequence $v_{1},\ldots,v_{k}=v$ of vertices such that $v_{1}\in\{w_{1},\ldots,w_{n}\}$, and $v_{i+1}$ is a vertex of the subgraph obtained from $v_{i}$ by sewing on the relator. That is, in $G$, we have
@@
For any $g\in G$, denote by $g_{u}$ the homomorphic image of $g$ under the canonical morphism $G\to\operatorname{FG}(u)$. Put $(w_{i})_{u}=u^{k_{i}}$ and let $K=\min{k_{i}}$.
Assume $u^{k}\in V(S\Gamma(w))$ for some $k\leq 0$, and put $W=\{w_{1},\ldots,w_{n}\}$. Then
@@
Notice that the only factors on the right hand side containing a negative power of $u$ are either in $W$, or are equal to $b^{-1}u^{-1}b$ or $a^{-1}u^{-1}a$. But no product containing $b^{-1}u^{-1}b$ or $a^{-1}u^{-1}a$ will freely reduce to $u^{k}$, as the only words in $\{a,ua,u,b,ub,b^{-1}u^{-1}b,a^{-1}u^{-1}a\}$ starting with $a^{-1}$ or $b^{-1}$ are $b^{-1}u^{-1}b$ and $a^{-1}u^{-1}a$. It follows that $k\geq K$.
@@
Denote the path metric on $S\Gamma(w)$ by $d$, and the word metric on $G$ with respect to the generating set $\{a,b,c,d\}$ by $d_{G}$. Define $v_{k}:=vbc^{-k}$ and $v_{k}^{\prime}:=vad^{-k}$. Observe that $v_{k}b^{-1}a=v_{k}^{\prime}$ in $G$, so $d_{G}(v_{k},v_{k}^{\prime})\leq 2$. We claim that $d(v_{k},v_{k}^{\prime})=2k+2$, which would imply the claim. Indeed, we shall show that any $v_{k}\to v_{k}^{\prime}$ path in $\Gamma(G)$ contains a vertex equal to $u^{k}$ for some $k$. It is then clear that the shortest such path contained in $S\Gamma(w)$ is the one through $v$ with length $22k+$ 2.
Recall that $G\cong\operatorname{FG}(a,b,u)$. Let $p$ be a path from $v_{k}\to v_{k}^{\prime}$ in $\Gamma(G)$ and let
$v_{k}=v_{k}g_{0},v_{k}g_{1},\ldots,v_{k}g_{n}=v_{k}^{\prime}$
be the vertices of $G$ traversed by $p$. Viewing $g_{i}$ as reduced words in $\operatorname{FG}(a,b,u)$, we in particular have $g_{0}=1$ and $g_{n}=v_{k}^{\prime}(v_{k})^{-1}=b^{-1}a$. Since
$g_{i+1}\in g_{i}$$\{a,b,c,d,a^{-1},b^{-1},c^{-1},d^{-1}\}^{\ast}=$
$=g_{i}$$\{a,b,b^{-1}ub,a^{-1}ua,a^{-1},b^{-1},b^{-1}u^{-1}b,a^{-1}u^{-1}a\}^{\ast}$
we claim that there must be some $i$ with $g_{i}=b^{-1}$. Indeed, consider the label $\ell(p)$ of $p$ as a word in $\{a,b,b^{-1}ub,a^{-1}ua,a^{-1},b^{-1},b^{-1}u^{-1}b,a^{-1}u^{-1}a\}^{\ast}$. By assumption, $\ell(p)$ freely reduces to $b^{-1}a$ – fix a reduction process and consider the last $b^{-1}$ in $\ell(p)$ left standing. Then in particular the prefix of $\ell(p)$ preceding this $b^{-1}$ freely reduces to $1$. Assume this $b^{-1}$ is a factor of $b^{-1}ub$ or $b^{-1}u^{-1}b$ in $\ell(p)$. Then the occurring $u$ or $u^{-1}$ would have to freely reduce, but this is impossible without the $b^{-1}$ reducing as well. It follows that the $b^{-1}$ remaining after the reduction process corresponds to a $b^{-1}$-edge traversed by $p$, let $g_{i}$ be the corresponding label. We then have $v_{k}g_{i}=vbc^{-k}g_{i}=u^{m}b(b^{-1}u^{-k}b)b^{-1}=u^{m-k}$ indeed. This completes the proof. ∎
@@
We have seen examples of unbounded group distortion; this naturally raises the question of when the distortion is bounded, either locally or uniformly. Our next aim is to explore the extent to which bounded distortion can be characterised as an algebraic property of the semigroup. Recall that an inverse semigroup is called an $F$-inverse semigroup if each $\sigma$-class has a greatest element with respect to the natural partial order; $F$-inverse semigroups are automatically $E$-unitary inverse monoids, with the identity element being the greatest element of the $\sigma$-class $E(S)$. In an $F$-inverse monoid we write $s^{\mathfrak{m}}$ for the greatest element in the $\sigma$-class of an element $s$. Notice that if $S$ is $F$-inverse and equipped with a proper length function, then $l_{S}(s)\leq l_{S}(s^{\mathfrak{m}})$ for every $s\in S$.
@@
It was observed in [20, Proposition 5.2.12] (in the language of coarse embeddings) that $F$-inverse monoids have uniformly bounded group distortion. In fact, the following result exactly characterises both locally and uniformly bounded group distortion as algebraic properties which can be seen as weak versions of being $F$-inverse.
@@
Let $G$ denote the group $S/\sigma$, and let $l_{S}$, $l_{G}$ be proper length functions on $S$ and $G$ respectively. First assume that every $\sigma$-class contains finitely many non-idempotent maximal elements, and that every ascending chain of non-idempotents is bounded above within the semigroup. For any $g\in G\setminus\{1\}$, let $m_{1}^{g},\ldots,m_{k_{g}}^{g}$ be the set of non-idempotent maximal elements of $S$ in the $\sigma$-class $g$. Since $G$ has bounded geometry we may define
@@
So the group distortion is uniformly bounded by $\phi$.
@@
[…]
@@
[…]
@@
This makes the class of special $F$-inverse monoids a very natural one to study from a geometric perspective. Our next aim is to characterize, in terms of the presentation, when an $E$-unitary special inverse monoid is $F$-inverse. We begin with an easy sufficient condition, which does not actually require that the inverse monoid is special.
@@
Let $g\in G$. By assumption $S\Gamma\cap gS\Gamma\neq\emptyset$, so the graph $S\Gamma\cup gS\Gamma$ is connected. Choose a $1\rightarrow g$ path in $S\Gamma\cup gS\Gamma$, and let $w_{g}$ be its label. We claim that $[w_{g}]_{S}$ is the greatest element of $\sigma^{-1}(g)$. Indeed, if $s\in S$ is such that $s{\sigma}=g$, then the embedded copy of $S\Gamma(s)$ inside $\Gamma(G)$ must contain $S\Gamma\cup gS\Gamma$, and therefore contains a path $ss^{-1}\xrightarrow{w_{g}}s$, proving that $[w_{g}]_{S}\geq s$. ∎
It turns out that when $S$ is special, $S\Gamma$ being sprawling can be described in terms of structural properties of the maximal group image $G$ and the prefix monoid $P$. Recall that a cancellative monoid is called right reversible if every pair of principal right ideals intersects. Right reversible monoids always give rise to a group of right quotients; we refer the reader to [3, Section 1.10] for details.
@@
Now assume (2). Then $P$ is right reversible and generates $G$ as a group. In particular any generator in $X$ can be expressed in terms of elements of $P$, and therefore any generator has to occur (in positive and/or negative form) in the relations, so (3) holds.
Conversely, assume (3). Since $P$ is right reversible, the set $PP^{-1}$ forms a subgroup in $G$ (see [3, Section 1.10]). To show that this subgroup equals $G$ it will suffice to show that $P$ generates $G$ as a group. Let $x\in X$ and suppose $w_{i}=pxq$ for some words $p,q\in(X\cup X^{-1})^{\ast}$ (the case when $w_{i}=px^{-1}q$ is similar). Then we have $p,px\in P$ and $[x]_{G}=[p^{-1}]_{G}[px]_{G}$, hence $P$ generates $G$ as required.
@@
Consider the special inverse monoid $S=\operatorname{Inv}\langle a,b\mid aba^{-n}b^{-1}=1\rangle$ for some fixed $n\geq 1$. This is $E$-unitary by [14] since the relator is cyclically reduced, and $G=Gp\langle a,b\mid aba^{-n}b^{-1}=1\rangle$ is the Baumslag-Solitar group $BS(1,n)$. It is not difficult to see that $P$ is the submonoid of $G$ generated by $a$ and $b$. Observe that $G=PP^{-1}$: for any $w\in\{a,b,a^{-1},b^{-1}\}^{*}$, we can apply reductions $a^{-1}b\to ba^{-n}$ and $b^{-1}a\to a^{n}b^{-1}$ along with group reductions to iteratively eliminate all subwords of the form $x^{-1}y$ where $x,y$ are positive letters. This results in a word $w^{\prime}$ with $[w^{\prime}]_{G}=[w]_{G}$ and $w^{\prime}=uv^{-1}$ for some $u,v\in\{a,b\}^{\ast}$, showing $[w]_{G}\in PP^{-1}$. It follows that $S$ is $F$-inverse.
One can show that right reversibility of the prefix monoid $P$ suffices for an $E$-unitary special inverse monoid to be $F$-inverse, even without the assumption that every generator appears in some relation. The proof relies on the fact that a free product of $F$-inverse monoids is always $F$-inverse. This latter fact can be deduced directly from the model of free products given in [15] as cactoid automata, but to avoid introducing technical definitions for which we have no further use here, we give a direct proof. All we use is that inverse monoids embed into their free products, and that the maximal group image of the free product of $S$ and $T$ is the group free product $S/\sigma\ast T/\sigma$.
@@
The free product of $F$-inverse monoids (in the category of inverse monoids) is $F$-inverse.

### v2

@@ 4. Distortion and (weak) $F$-Inverse properties
To prove that $S/\sigma=G$, observe that the group defined by regarding the presentation above as a group presentation is indeed $G$, as (R1) ensures $e]_{G}=1$, (R2) yields a presentation for $G$ by definition, and the relations in (R3) are then automatic. The fact that $S$ is $E$-unitary follows immediately from the relations in (R1) and (R2).
Now let $\Gamma$ denote the Cayley graph of $G$ with respect to the generating set $X\cup\{e\}$. Clearly $\Gamma$ is obtained from the Cayley graph of $G$ with respect to $X$ by adding an $e$-labelled loop at each vertex. Let $\Omega$ be the subgraph of $\Gamma$ consisting of $\Delta$ with an extra $e$-labelled loop at $1$. We claim that $\Omega$ satisfies conditions of Lemma 2.3 to be isomorphic to the Schützenberger graph of $e$. Indeed, it clearly contains the path (in other words, the loop) labelled $e$ starting at $1$. It remains to show that if one side of a defining relation of $S$ labels a path between two vertices in $\Omega$, then the other side of the relation labels a path between the same two vertices. For the relations of type (R1) and (R2) notice that if $w=e$ or $w\in(X\cup X^{-1})^{\ast}$ with $w]_{G}=1$ then every path in $\Gamma$ labelled $w$ is closed; clearly this implies that $w$ labels a path between two vertices if and only if $w^{2}$ does so.
@@
The inverse monoid $S=\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}c^{-1}cd^{-1}d=1\rangle$ is an $E$-unitary inverse monoid without locally bounded group distortion. Indeed, every Schützenberger graph of $S$ embeds in the Cayley graph of the maximum group image with unbounded distortion.
@@
First, observe that the inverse monoid $S$ is the same as the one defined by the presentation $\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1,c^{-1}c=1,d^{-1}d=1\rangle$. This is an idempotent-pure quotient of the one-relator inverse monoid $\operatorname{Inv}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1\rangle$. The latter is $E$-unitary by [14] since the relator is a cyclically reduced word, and a quotient of an $E$-unitary inverse monoid by a congruence contained in the minimum group congruence is easily seen to be itself $E$-unitary, so $S$ itself is $E$-unitary.
The maximal group image of $S$ is $G=\operatorname{Gp}\langle a,b,c,d\mid bcb^{-1}ad^{-1}a^{-1}=1\rangle$. If we introduce a new generator $u=bcb^{-1}$, then
@@
For convenience, we will identify the elements of $G$ with freely reduced words over the generators $a$, $b$, $u$ and their inverses. In particular, in the Cayley graph $\Gamma(G)$ of $G$ with respect to the generating set $a,b,c,d$, we identify vertices of $\Gamma(G)$ with $\operatorname{FG}(a,b,u)$, while the edges are labelled by the set $\{a,b,c,d\}$. It is helpful to keep in mind this picture depicting the relations in $G$:
ccdduuaabbaabb
The Cayley graph $\Gamma(G)$ is obtained from the Cayley graph of $\operatorname{FG}(a,b,u)$ (with respect to the generating set $a,b,u$) by replacing every $u$-edge with the surrounding rectangle and determinising. In particular, observe that the black edges of the graph below form a subgraph of $\Gamma(G)$:
@@
ccddaabbaabbccdd
In particular, $S\Gamma(w)$ contains the graph below as a subgraph at every vertex $v$.
@@
We furthermore claim that there exists some vertex $v$ of $S\Gamma(w)$ such that no path with label of the form $bc^{-k}b^{-1}$, or $ad^{-k}a^{-1}$ $(k\in\mathbb{N})$ is readable from $v$ in $S\Gamma(w)$. Note that $bc^{-k}b^{-1}]_{G}=[ad^{-k}a^{-1}]_{G}=u^{-k}$, so it suffices to show that there exists some vertex $v$ of $\Gamma$ such that $vu^{-k}\notin V(S\Gamma(w))$ for any positive $k$.
Let $w_{1},\ldots,w_{n}\in G$ be the vertices traversed by $w$ and define $W=\{w_{1},\dots w_{n}\}$. Then for any $v\in V(S\Gamma(w))$, there is a sequence $v_{1},\ldots,v_{k}=v$ of vertices such that $v_{1}\in W$, and $v_{i+1}$ is a vertex of the subgraph of $\Gamma_{G}$ obtained by sewing on the relator at $v_{i}$. That is, in $G$, we have
@@
Let $k_{i}$ denote the number of occurrences of $u^{-1}$ in $w_{i}$ (viewed as a reduced word in $\operatorname{FG}(a,b,u)$) and let $K=-\max{k_{i}}$.
Assume $u^{k}\in V(S\Gamma(w))$ for some $k\leq 0$, we claim that then $k\geq K$. Then we have
@@
[…]
@@
Denote the path metric on $S\Gamma(w)$ by $d$, and the word metric on $G$ with respect to the generating set $\{a,b,c,d\}$ by $d_{G}$. For $k\geq 1$ define $u_{k}:=[vbc^{-k}]_{G}=u^{m-k}b$ and $u_{k}^{\prime}:=[vad^{-k}]_{G}=u^{m-k}a$; it is easily seen that the elements $vbc^{-k}$ and $vad^{-k}$ are right-invertible in $S$ and it follows that $u_{k}$ and $u_{k}^{\prime}$ are vertices of $S\Gamma(w)$. Observe that $u_{k}b^{-1}a=u_{k}^{\prime}$, so $d_{G}(u_{k},u_{k}^{\prime})\leq 2$. We claim that $d(u_{k},u_{k}^{\prime})=2k+2$, which would imply the that the embedding of $S\Gamma(w)$ has unbounded distortion. Indeed, we shall show that every $u_{k}\to u_{k}^{\prime}$ path in $\Gamma(G)$ contains a vertex $u^{n}$ for some $n$. It is then clear that the shortest such path contained in $S\Gamma(w)$ is the one through $v$ with length $2k+2$.
For any path $p$ in $\Gamma(G)$, let $\ell(p)$ denote the label of $p$ as a word in the alphabet $\{a,b,c,d\}^{\pm 1}$, and let $\ell_{u}(p)$ be the word obtained from this label by replacing any occurrences $c^{\pm 1}$ and $d^{\pm 1}$ with $b^{-1}u^{\pm 1}b$ and $a^{-1}u^{\pm 1}a$ respectively, that is,
$\ell_{u}(p)\in\{a,b,b^{-1}ub,a^{-1}ua,a^{-1},b^{-1},b^{-1}u^{-1}b,a^{-1}u^{-1}a\}^{\ast}.$
Now suppose $p$ is a path from $v_{k}\to v_{k}^{\prime}$ in $\Gamma(G)$; then $\ell_{u}(p)$ must freely reduce to $b^{-1}a$. Furthermore, $p$ traverses a vertex $u^{n}\in\Gamma(G)$ if and only if $p$ has a prefix $p^{\prime}$ such that $\ell_{u}(p^{\prime})$ freely reduces to $v_{k}^{-1}u^{n}=b^{-1}u^{k-m+n}$, thus our goal is to show that such prefix exists. (Notice that not all prefixes of $\ell_{u}(p)$ come from prefixes of $\ell(p)$, or equivalently of $p$, so this claim is stronger than simply requiring $\ell_{u}(p)$ to have a prefix freely reducing to $b^{-1}u^{k-m+n}$.)
We prove the following stronger statement: if $q$ be a path in $\Gamma(G)$ such that $\ell_{u}(q)$ freely reduces to a word of the form $b^{-1}u^{l}a$, then $q$ has a prefix $q^{\prime}$ such that $\ell_{u}(q^{\prime})$ freely reduces to a word of the form $b^{-1}u^{n}$. Note that the latter condition is equivalent to requiring that $q$ has a suffix $q^{\prime\prime}$ with $\ell_{u}(q^{\prime\prime})$ freely reducing to a word of the form $u^{n^{\prime}}a$. We prove the claim by induction on the length of the path $q$. The statement is obviously true when $\ell(q)=b^{-1}a$, which is the shortest path satisfying the conditions. Now let $q$ be a path such that $\ell_{u}(q)$ freely reduces to $b^{-1}u^{l}a$ for some $l\geq 1$ and assume the statement is true for all shorter paths of this form.
Fix some reduction sequence of $\ell_{u}(q)$ to a freely reduced word. Then we may factorise $\ell_{u}(q)$ as $w_{1}\cdot b^{-1}\cdot w_{2}\cdot a\cdot w_{3}$ where $w_{1}$ and $w_{3}$ freely reduce to $1$ and $w_{2}$ freely reduces to $u^{l}$. Consider the prefix $w_{1}b^{-1}$. If this prefix is the label of some prefix $q^{\prime}$ of the path $q$ then since it freely reduces to $b^{-1}$ we are done. Dually, if the suffix $aw_{3}$ is the label of some suffix $q^{\prime\prime}$ then the complementary prefix $q^{\prime}$ has label freely reducing to $b^{-1}u^{l}$ and we are done. Otherwise, the $b^{-1}$ ending the prefix $w_{1}b^{-1}$ and the $a$ beginning the suffix $aw_{3}$ must come respectively from edges in $q$ labelled $b^{-1}u^{\varepsilon_{1}}b$ and $a^{-1}u^{\varepsilon_{2}}a$ for some $\varepsilon_{1},\varepsilon_{2}\in\{1,-1\}$. In this case
$\ell_{u}(q)=w_{1}\cdot b^{-1}\underbrace{u^{\varepsilon_{1}}b\cdot\widehat{w}\cdot a^{-1}u^{\varepsilon_{2}}}_{w_{2}}a\cdot w_{3}.$
Since $w_{2}$ freely reduces to $u^{l}$, $\widehat{w}$ must freely reduce to $b^{-1}u^{l-\varepsilon_{1}-\varepsilon_{2}}a$. Observe that $\widehat{w}$ is the label of a subpath of $q$ which satisfies the conditions of the claim and is shorter than $q$. So by the inductive hypothesis there exists a prefix $\widehat{q}^{\prime}$ of $\widehat{q}$ such that $\ell_{u}(\widehat{q}^{\prime})$ freely reduces to a word of the form $b^{-1}u^{n}$. Let now $q^{\prime}$ be the prefix of $q$ with
$\ell_{u}(q^{\prime})=w_{1}\cdot b^{-1}u^{\varepsilon_{1}}b\cdot\ell_{u}(\widehat{q}^{\prime}),$
which we can reduce freely as $w_{1}\cdot b^{-1}u^{\varepsilon_{1}}b\cdot\ell_{u}(\widehat{q}^{\prime})\to b^{-1}u^{\varepsilon_{1}}b\cdot b^{-1}u^{n}\to b^{-1}u^{\varepsilon_{1}+n}$, proving the claim. Applying the claim to the path $p$ completes the proof of the theorem. ∎
@@
We have seen examples of unbounded group distortion; this naturally raises the question of when the distortion is bounded, either locally or uniformly. Our next aim is to explore the extent to which bounded distortion can be characterised as an algebraic property of the semigroup, related to being $F$-inverse. In an $F$-inverse monoid we write $s^{\mathfrak{m}}$ for the greatest element in the $\sigma$-class of an element $s$. Notice that if $S$ is $F$-inverse and equipped with a proper length function, then $l_{S}(s)\leq l_{S}(s^{\mathfrak{m}})$ for every $s\in S$.
@@
It was observed in [19, Proposition 5.2.12] (in the language of coarse embeddings) that $F$-inverse monoids have uniformly bounded group distortion. In fact, the following result exactly characterises both locally and uniformly bounded group distortion as algebraic properties which can be seen as weak versions of being $F$-inverse.
@@
Let $G$ denote the group $S/\sigma$, and let $l_{S}$, $l_{G}$ be proper length functions on $S$ and $G$ respectively. First assume that every $\sigma$-class contains finitely many non-idempotent maximal elements, and that every ascending chain of non-idempotents is bounded above within the semigroup. For each $g\in G\setminus\{1\}$, let $m_{1}^{g},\ldots,m_{k_{g}}^{g}$ be the non-idempotent maximal elements of $S$ in the $\sigma$-class $g$. Since $G$ has bounded geometry we may define
@@
So by Proposition 4.4, the group distortion is uniformly bounded by $\phi$.
@@
Denote the determinisation map on the vertices by $f\colon V(S\Gamma(s)\vee S\Gamma(t^{-1}))\to V(S\Gamma(st^{-1}))$. Our aim is to describe when two vertices are identified by $f$. Consider the relation $\sim$ on $V(S\Gamma(s)\vee S\Gamma(t^{-1}))$, where for $u\in V(S\Gamma(s)),v\in V(S\Gamma(t^{-1}))$, we define $u\sim v$ if there are is a path $z\xrightarrow{w}u$ in $S\Gamma(s)$, and $z\xrightarrow{w}v$ in $S\Gamma(t)$ for some word $w\in(X\cup X^{-1})^{\ast}$, and take the symmetric and reflexive closure.
@@
It remains to show that $S\Gamma(s)\vee S\Gamma(t^{-1})/\sim$ is deterministic. Let $u_{1},v_{1},u_{2},v_{2}$ be vertices of $S\Gamma(s)\vee S\Gamma(t^{-1})$ with edges $u_{1}\xrightarrow{x}v_{1}$ and $u_{2}\xrightarrow{x}v_{2}$ such that $u_{1}\sim u_{2}$. We need to show that $v_{1}\sim v_{2}$. If $u_{1}=u_{2}$, then we must either have $v_{1}=v_{2}$, or $u_{1}=u_{2}=z$ and $v_{1}\in V(S\Gamma(s))$, $v_{1}\in V(S\Gamma(t^{-1}))$, or vice versa. In either case, we must have $v_{1}\sim v_{2}$ by definition. Otherwise, without loss of generality, we may assume that $u_{1}\in V(S\Gamma(s))$, $u_{2}\in V(S\Gamma(t^{-1}))$, and there is a word $w\in(X\cup X^{-1})^{\ast}$ with paths $z\xrightarrow{w}u_{1}$ and $z\xrightarrow{w}u_{2}$ lying in $S\Gamma(s)$ and $S\Gamma(t^{-1})$ respectively. If $u_{1}=z$, then, because $S$ is $E$-unitary and $z\xrightarrow{w}u_{1}$ is a path in a Schützenberger graph, we must have that $w_{S}$ is idempotent, and $z\xrightarrow{w}u_{2}$ is also a cycle, making $u_{1}=u_{2}$ – a contradiction. So we must have $u_{1}\neq z$, and similarly $u_{2}\neq z$. It follows that $u_{1}\xrightarrow{x}v_{1}$ is contained in $S\Gamma(s)$ and $u_{2}\xrightarrow{x}v_{2}$ contained in $S\Gamma(t^{-1})$. But then there are paths $z\xrightarrow{wx}v_{1}$ in $S\Gamma(s)$ and $z\xrightarrow{wx}v_{2}$ in $S\Gamma(t^{-1})$, so $v_{1}\sim v_{2}$ indeed. This completes the proof that $S\Gamma(s)\vee S\Gamma(t^{-1})/\sim$ is deterministic, establishing the claim that $\sim$ coincides with $\ker f$.
Since $s\mathrel{\sigma}t$ and $S$ is $E$-unitary, $st^{-1}$ must be idempotent, in particular, $f$ identifies the vertex $ss^{-1}$ of $\Gamma(s)$ and the vertex $t$ of $S\Gamma(t^{-1})$, so there must be paths $z\xrightarrow{w}ss^{-1}$ in $S\Gamma(s)$, and $z\xrightarrow{w}t^{-1}$ in $S\Gamma(t^{-1})$ for some word $w\in(X\cup X^{-1})^{\ast}$. In particular, $w^{-1}$ labels a path in $S\Gamma(s)$ from $ss^{-1}$ to $s$ and hence $w^{-1}]_{S}=[w]_{S}^{-1}\geq s$, and similarly $w$ labels a path in $S\Gamma(t^{-1})$ from $tt^{-1}$ to $t^{-1}$ and hence $w]_{S}\geq t^{-1}$ and so $w]_{S}^{-1}\geq t$. This shows that $w]_{S}^{-1}$ is a common upper bound for $s$ and $t$ as required. ∎
@@
This makes the class of special $F$-inverse monoids a very natural one to study from a geometric perspective. Our next aim is to characterise, in terms of the presentation, when an $E$-unitary special inverse monoid is $F$-inverse. We begin with an easy sufficient condition, which does not actually require that the inverse monoid is special.
@@
Let $g\in G$. By assumption $S\Gamma\cap gS\Gamma\neq\emptyset$, so the graph $S\Gamma\cup gS\Gamma$ is connected. Choose a $1\rightarrow g$ path in $S\Gamma\cup gS\Gamma$, and let $w_{g}$ be its label. We claim that $w_{g}]_{S}$ is the greatest element of $\sigma^{-1}(g)$. Indeed, if $s\in S$ is such that $s{\sigma}=g$, then the embedded copy of $S\Gamma(s)$ inside $\Gamma(G)$ must contain $S\Gamma\cup gS\Gamma$, and therefore contains a path $ss^{-1}\xrightarrow{w_{g}}s$, proving that $w_{g}]_{S}\geq s$. ∎
It turns out that when $S$ is special, $S\Gamma$ being sprawling can be described in terms of structural properties of the maximal group image $G$ and the prefix monoid $P$. Recall that a cancellative monoid is called right reversible if every pair of principal right ideals intersects. If $P$ is a right reversible monoid, then it always give rise to a group of right quotients, that is, a (necessarily unique) group $G$ containing $P$ as a submonoid where every element can be expressed in the form $pq^{-1}$ with $p,q\in P$. Conversely, every cancellative monoid with a group of right quotients is right reversible. We refer the reader to [3, Section 1.10] for more detail.
@@
Now assume (2). Then $P$ is right reversible and generates $G$ as a group. In particular every generator in $X$ can be expressed in terms of elements of $P$, and therefore every generator has to occur (in positive and/or negative form) in the relations, so (3) holds.
Conversely, assume (3). Since $P$ is right reversible, the set $PP^{-1}$ forms a subgroup in $G$ (see [3, Section 1.10]). To show that this subgroup equals $G$ it will suffice to show that $P$ generates $G$ as a group. Let $x\in X$ and suppose $w_{i}=pxq$ for some words $p,q\in(X\cup X^{-1})^{\ast}$ (the case when $w_{i}=px^{-1}q$ is similar). Then we have $p,px\in P$ and $x]_{G}=[p^{-1}]_{G}[px]_{G}$, hence $P$ generates $G$ as required.
@@
Consider the special inverse monoid $S=\operatorname{Inv}\langle a,b\mid aba^{-n}b^{-1}=1\rangle$ for some fixed $n\geq 1$. This is $E$-unitary by [14] since the relator is cyclically reduced, and $G=Gp\langle a,b\mid aba^{-n}b^{-1}=1\rangle$ is the Baumslag-Solitar group $BS(1,n)$. It is not difficult to see that $P$ is the submonoid of $G$ generated by $a$ and $b$. Observe that $G=PP^{-1}$: for any $w\in\{a,b,a^{-1},b^{-1}\}^{*}$, we can apply reductions $a^{-1}b\to ba^{-n}$ and $b^{-1}a\to a^{n}b^{-1}$ along with group reductions to iteratively eliminate all subwords of the form $x^{-1}y$ where $x,y$ are positive letters. This results in a word $w^{\prime}$ with $w^{\prime}]_{G}=[w]_{G}$ and $w^{\prime}=uv^{-1}$ for some $u,v\in\{a,b\}^{\ast}$, showing $w]_{G}\in PP^{-1}$. It follows that $S$ is $F$-inverse.
We shall need the following elementary fact about free products of $F$-inverse monoids, which may well be known but which we have been unable to find stated in the literature. The proof uses $F$-inverse covers. Given an inverse semigroup $S$, and $F$-inverse cover of $S$ is an $F$-inverse monoid $F$ with an idempotent separating surjective morphism $F\to S$.
@@
The free product of two $F$-inverse monoids (in the category of inverse monoids) is $F$-inverse.